OpenStudy (anonymous):

what is an exponent with a prime number as a base and it equals 50?

OpenStudy (saifoo.khan):

1.

OpenStudy (saifoo.khan):

\[50^1 = 50\]

OpenStudy (anonymous):

1 is not prime

OpenStudy (saifoo.khan):

hmm, right.

OpenStudy (anonymous):

i really need help on my homework

OpenStudy (anonymous):

what is a exponent with a prime number as a base and equals 50?

OpenStudy (anonymous):

so what needs to be prime the number(x) or the exponent(e)? \[x ^{e}\]

OpenStudy (anonymous):

the base of the exponent needs to be prime

OpenStudy (anonymous):

help

OpenStudy (anonymous):

so the x in the above equation?

OpenStudy (saifoo.khan):

\[2^1 \ \times 25\]

OpenStudy (anonymous):

ok thanks.

OpenStudy (anonymous):

wait does that equal 50?

OpenStudy (anonymous):

its 2 *25 which does equal 25 but I dont think its what you are looking for

OpenStudy (anonymous):

it is not i needs to equal 50.

OpenStudy (anonymous):

corey can you help?

OpenStudy (anonymous):

does the exponent have to be a whole number?

OpenStudy (anonymous):

yes if you have a whole number exponent that would be great but it does not have to be.

OpenStudy (anonymous):

the replys take forever

OpenStudy (anonymous):

because you could set it up like so: \[2^{x} = 50\] because 2 is prime and 50 is what you need you would solve for x which is ln(50)/ln(2) = 5.6438

OpenStudy (anonymous):

no but i need to know what x is

OpenStudy (anonymous):

x is 5.6438 which would give you 49.99 rounded up is 50

OpenStudy (anonymous):

cant be rounded

OpenStudy (anonymous):

ok x would be ln(50)/ln(2) which gives you 50 exactly when the base is 2

OpenStudy (anonymous):

so what would the answer look like?

OpenStudy (anonymous):

\[2^{\ln(50)/\ln(2)}\]

OpenStudy (anonymous):

is that an exponent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aww man i have to go eat hurry type

OpenStudy (anonymous):

you could set it up for many primes that way, but the exponent would change